OpenStudy (anonymous):

Evaluate sin 60° without using a calculator by using ratios in a reference triangle.

OpenStudy (anonymous):

Answer: sqrt(3)/2.

OpenStudy (anonymous):

But how do you get that answer?

OpenStudy (anonymous):

Just memorize it. This will occur very often. Sin 30=1/2, Sin 60=sqrt(3)/2, Cos 30=sqrt(3)/2, Cos 60=1/2.

OpenStudy (anonymous):

So when I type my answer and explain why I just reference the unit circle?

OpenStudy (anonymous):

You'll learn the reference angle later. For now, just memorize the basic trigonometric values like this one.

OpenStudy (anonymous):

okay thank you(:

OpenStudy (anonymous):

No problem.